Logiweb diagnose of prove

Up Help

No errors found

The pyk compiler, version 0.grue.20060417+ by Klaus Grue GRD-2006-12-07.UTC:17:24:53.972444 = MJD-54076.TAI:17:25:26.972444 = LGT-4672229126972444e-6